OpenStudy (anonymous):

Identify the zeros of the function: y=(x+3)(x+12)

OpenStudy (angela210793):

x=-3 or x=-12

OpenStudy (anonymous):

-3, -12

OpenStudy (anonymous):

Explained: x+3 = 0 x = 0 - 3 x = -3 ---------- x + 12 = 0 x = 0 - 12 x = -12